# Pragmatics in the Era of Large Language Models: A Survey on Datasets, Evaluation, Opportunities and Challenges

### v1

@@ Abstract
Understanding pragmatics—the use of language in context—is crucial for developing NLP systems capable of interpreting nuanced language use. Despite recent advances in language technologies, including large language models, evaluating their ability to handle pragmatic phenomena such as implicatures and references remains challenging. To advance pragmatic abilities in models, it is essential to understand current evaluation trends and identify existing limitations. In this survey, we provide a comprehensive review of resources designed for evaluating pragmatic capabilities in NLP, categorizing datasets by the pragmatics phenomena they address. We analyze task designs, data collection methods, evaluation approaches, and their relevance to real-world applications. By examining these resources in the context of modern language models, we highlight emerging trends, challenges, and gaps in existing benchmarks. Our survey aims to clarify the landscape of pragmatic evaluation and guide the development of more comprehensive and targeted benchmarks, ultimately contributing to more nuanced and context-aware NLP models.
@@ A Survey on Datasets, Evaluation, Opportunities and Challenges
Bolei Ma1, Yuting Li2, Wei Zhou3,4, Ziwei Gong5, Yang Janet Liu1, Katja Jasinskaja2, Annemarie Friedrich3, Julia Hirschberg5, Frauke Kreuter1,6, Barbara Plank1 1LMU Munich & Munich Center for Machine Learning, 2University of Cologne, 3University of Augsburg, 4Bosch Center for Artificial Intelligence, 5Columbia University, 6University of Maryland, College Park Equal contributions. bolei.ma@lmu.de, yuting.li@uni-koeln.de, wei.zhou@de.bosch.com, zg2272@columbia.edu
@@ 1 Introduction
In linguistics, pragmatics studies how context influences the meaning of language Huang (2017); Xiang et al. (2024); Birner (2012), and how people use language in real-life situations to convey implied meanings, emotions and intentions. Foundational work in this field, e.g. Grice (1975)’s work on implicature and the cooperative principle, Austin (1975)’s idea of speech acts, and Sperber and Wilson (1986)’s exploration of contextual inference laid the groundwork of the study of language use. These concepts continually influence linguistic studies and also provide insights for computational methods Mann (1980); Saygin and Cicekli (2002); Hovy and Yang (2021); Cambria (2024).
@@
However, traditional Natural Language Processing (NLP) models, while competent in syntactic parsing and semantic analysis, struggle with meaning that extends beyond the literal definition of words. This gap is where pragmatics becomes essential. Early approaches, like rule-based systems, use explicitly coded rules for knowledge representation and match input with a knowledge base for response generation Grosan et al. (2011); Bajwa and Choudhary (2006). Later, statistical models applied probability theory to predict language behavior Johnson (2009), while deep learning, with transformer-based models like BERT Devlin et al. (2019) and GPT Brown et al. (2020), transformed NLP by enabling the generation of contextually relevant text. Yet their performance in understanding pragmatics remains limited Hu et al. (2023); Sileo et al. (2022); Park et al. (2024b).
@@
Recently, the emergence of Large Language Models (LLMs) has intensified the need to evaluate their human-like communication abilities, particularly for nuanced use cases requiring sophisticated pragmatic reasoning Hu et al. (2023); Ruis et al. (2023); Yerukola et al. (2024). As LLMs are increasingly deployed in real-world applications, validating their ability to understand and generate contextually appropriate and pragmatically accurate responses is crucial to ensure effective and trustworthy human-computer interactions. While recent advances in LLMs have demonstrated impressive capabilities in generating coherent and contextually appropriate text, their pragmatic competence remains inadequately evaluated Hu et al. (2023); Kwon et al. (2023); Chang et al. (2024).
@@
Important questions arise: What resources have been developed to evaluate the pragmatic capabilities of NLP models, and, more importantly, how can we leverage pragmatics to guide the advancement of LLMs, formalize a comprehensive framework for evaluating pragmatics, and further advance the study of pragmatics in linguistics? To answer these questions, we conduct a comprehensive survey of the resources available for evaluating pragmatics in NLP: In §3 we review how pragmatics phenomena have been evaluated across various NLP tasks. In §4 we summarize how previous works build a pragmatics dataset. In §5 we present the metrics and evaluation techniques used in current research. In §6 we highlight gaps and offer recommendations for future research.
@@
Figure 1: A taxonomy of pragmatics phenomena.
@@ 2 Core Concepts in Pragmatics
2.1 Pragmatics Phenomena in Linguistics
@@
Context is the foundation of interpretating pragmatics, encompassing both linguistic context (e.g., surrounding clauses or sentences) and extra-linguistic context, including temporal, spatial, and social factors Huang (2017).
Deixis.
@@
In particular, deictic expressions depend heavily on the situational context of the utterance for their meaning. These expressions, which include personal deixis (e.g., “I”, “you”), spatial deixis (e.g., “this”, “that”), and temporal deixis (e.g., “now”, “then”), are interpreted based on factors such as the speaker’s location, the time of the utterance, and the prior discourse Levinson (2006); Stapleton (2017); Levinson (1983).
Implicature.
@@
Grice (1975)’s notion of implicature describes how speakers imply additional meanings without stating them explicitly. For example, if Ann asks, “Do you sell paste?” and Bill replies, “I sell rubber cement”, the implication is that Bill does not sell paste Hirschberg (1985). Implicature results from enriching the context with additional assumptions to comply with the cooperative principle and the conversational maxims: truthfulness (Maxim of Quality), informativeness (Maxim of Quantity), relevance (Maxim of Relevance), and clarity (Maxim of Manner) (Grice, 1975).
Presupposition.
@@
Presuppositions are the implicit assumptions that must hold true for an utterance to be meaningful and understood Stalnaker et al. (1977). For instance, the sentence “The Queen of England is bald” presupposes that England has a unique Queen, even though this fact is not directly stated Stalnaker et al. (1977).
Speech Acts and Intent Recognition.
@@
Understanding speech acts is critical for understanding speaker intent. A speech act is an utterance that not only conveys information but also performs an action Austin (1975). Each utterance can be classified as performing an act that fulfills a communicative purpose, such as such as assertion, suggestion or description. For example, the utterance “Can you open the window?” might literally ask about the addressee’s ability or, depending on the context, function as an indirect request.
Discourse and Coherence.
@@
In linguistics, discourse refers to language use beyond sentence level Guardado (2018). Discourse structure is studied by examining how sentence and larger text units are interconnected through coherence relations such as elaboration, explanation and contrast. Some theoretical approaches offer hierarchical models that map these relations to reveal the underlying structure of texts (for a detailed overview of Rhetorical Structure Theory, see Mann and Thompson, 1988), while other frameworks provide a formal semantic account that integrates these coherence relations into dynamic discourse representation (for further insights into Segmented Discourse Representation Theory, see Asher and Lascarides, 2003).
Social Pragmatics.
@@
Social pragmatics broadens traditional theories by exploring how social factors like culture, power, gender, and interpersonal dynamics influence language use. Brown and Levinson (1987) introduced the concept of “face” (self-esteem), showing how politeness strategies manage relationships and preserve social harmony. Gender differences in communication, as examined by Tannen (1990) and Eckert and McConnell-Ginet (1992), highlight contrasting conversational styles between men and women. Cultural variations further complicate communication, as Scollon et al. (2011) demonstrate that politeness norms differ across cultures. In the digital era, social pragmatics also informs computer-mediated communication, where cues like emoticons convey emotions and tone Herring et al. (2013).
2.2 Pragmatics Phenomena in the Survey
@@
We next examine the pragmatics phenomena addressed in NLP, which were studied in the surveyed works. An overview of the survey methods and inclusivity is presented in Appendix §A. Since pragmatic concepts are interconnected, there are overlaps between categories. Therefore, we focus on the key aspects that each paper addresses. Figure 1 gives an overview of the phenomena and papers.
Context and Deixis.
@@
Context and deixis in NLP evaluation tasks usually assess a model’s ability to interpret inputs based on the situational or linguistic context. The examination of the context and deixis is the basis of evaluating the models’ pragmatic ability as the models rely on the context to respond, similar to the human response process Greasley and Owen (2016); van Dijk et al. (2023); Ma et al. (2024a). Datasets or frameworks like Min et al. (2020); Sravanthi et al. (2024); Li et al. (2023a); Qi et al. (2023); Sileo et al. (2022); Hu et al. (2023); Shaikh et al. (2023); Park et al. (2024b, a); de Vries et al. (2017) explicitly address deixis as one of their pragmatic phenomena, requiring models to resolve references in context. Additionally, Li et al. (2023a) and Qi et al. (2023) incorporate context-dependent reasoning, requiring models to interpret utterances within multi-turn dialogues. Monroe et al. (2017, 2018) explore pragmatic language use in color reference tasks, demonstrating how context influences referential choices. Westera et al. (2020b) and Zeyrek et al. (2018) also emphasize the role of context in discussion structure and question-under-discussion frameworks. These datasets highlight the role of context in grounding language understanding.
Implicature and Presupposition.
@@
A recurring theme across many papers is the study of implicature, i.e., testing the model’s inference ability based on the textual input across literal meaning. It explores how meaning is implied rather than explicitly stated, requiring models to infer intentions and assumptions beyond literal interpretations. Different types of implicatures have been studied: Qi et al. (2023), Hu et al. (2023) and Halat and Atlamaz (2024) focus on conversational implicatures, testing models’ ability to infer pragmatic meanings in dialogue. Nizamani et al. (2024) investigate scalar implicatures, particularly with gradable adjectives. Zheng et al. (2021) use a grammar-based approach to generate dialogues with intricate implicatures. Presupposition is another aspect of implicature and influences the model’s inference ability on the implications. Louis et al. (2020); Damgaard et al. (2021); Müller and Plank (2024) focus on the indirectness in dialogue. Sap et al. (2020) work on the implicatures posted in social media posts and study the social biases. Kameswari et al. (2020) use pragmatic presupposition to enhance bias detection in political news. Moreover, additional works Koyano et al. (2022); Srikanth et al. (2024); Kim et al. (2024) use pragmatic inferences to understand the meanings of the given input.
Speech Acts and Intent Recognition.
@@
Speech acts have been explicitly studied in the surveyed works, usually in cases where there are requests, commands, or promises, given to the model and test the model’s intent recognition. Li et al. (2023a) include tasks for pragmatic identification and reasoning, requiring models to recognize speech acts in multi-turn dialogues. Reinig et al. (2024) provide a fine-grained annotation of speech acts in parliamentary debates. Braga et al. (2006) link prosodic patterns to speech acts and discourse events. A few papers have explicitly different “speakers” involved in the evaluation frameworks and assess how well the models play as the “speakers” in the dialogues, performing the commended task. Khani et al. (2018) explore how players infer intentions and generate pragmatic messages in collaborative tasks, focusing on the rational speech act between the two agents (the “speaker” and the “listener”). Shaikh et al. (2023) involve a multi-turn collaborative two-player game based on the codenames. Ollagnier (2024) also incorporates speech act-like annotations, such as “attack” and “defend”, to capture the pragmatic roles of messages in multiparty chats. Both Zhang et al. (2018) and Welleck et al. (2019) talk about the personas, who are the agents making requests, in dialogue and work on the consistency between persona and dialogue.
Discourse and Coherence.
@@
Discourse and coherence has been studied in NLP with a focus on parsing text into discourse relations (e.g., Miltsakaki et al., 2004; Prasad et al., 2008, 2018) and on modeling dimensions of coherence (e.g., Lai and Tetreault, 2018; Wu et al., 2023). In addition, discourse structures play a crucial role in dialogue and communication. Westera et al. (2020b) and Zeyrek et al. (2018) investigate the role of evoked questions in discourse structure using TED talks. Asher et al. (2016) explore discourse structure in multi-party dialogues and Reinig et al. (2024) study multi-party dialogues in German parliament. Moreover, discourse modes and aspect selection are influenced by contextual and pragmatic factors, connecting coherence to pragmatic interpretation (Mavridou et al., 2015). For a comprehensive overview of situation types and aspects, see Friedrich et al. (2023).
Social Pragmatics.
@@
Social pragmatics explores how language use is shaped by social norms, power dynamics, and cultural contexts, including the roles of agents in communication. Ollagnier (2024) introduces a tagset for annotating discursive roles in cyberbullying, while Sap et al. (2019) and Sap et al. (2020) provide benchmarks for reasoning about social norms, biases, and power dynamics. Shaikh et al. (2023) highlights the role of shared cultural knowledge in communication through a collaborative word game, and Zhang et al. (2018) and Welleck et al. (2019) emphasize using personal and social context to improve dialogue systems. Additionally, Yang et al. (2021) examine how pragmatic features vary between neurotypical and neurodiverse speakers, particularly adults with autism. These datasets collectively demonstrate the intersection of social pragmatics with other phenomena like discourse, implicature, and speech acts.

### v3

@@ Abstract
Understanding pragmatics—the use of language in context—is crucial for developing NLP systems capable of interpreting nuanced language use. Despite recent advances in language technologies, including large language models, evaluating their ability to handle pragmatic phenomena such as implicatures and references remains challenging. To advance pragmatic abilities in models, it is essential to understand current evaluation trends and identify existing limitations. In this survey, we provide a comprehensive review of resources designed for evaluating pragmatic capabilities in NLP, categorizing datasets by the pragmatic phenomena they address. We analyze task designs, data collection methods, evaluation approaches, and their relevance to real-world applications. By examining these resources in the context of modern language models, we highlight emerging trends, challenges, and gaps in existing benchmarks. Our survey aims to clarify the landscape of pragmatic evaluation and guide the development of more comprehensive and targeted benchmarks, ultimately contributing to more nuanced and context-aware NLP models.
@@ A Survey on Datasets, Evaluation, Opportunities and Challenges
Bolei Ma1, Yuting Li2, Wei Zhou3,4, Ziwei Gong5, Yang Janet Liu1, Katja Jasinskaja2, Annemarie Friedrich3, Julia Hirschberg5, Frauke Kreuter1,6, Barbara Plank1 1LMU Munich & Munich Center for Machine Learning, 2University of Cologne, 3University of Augsburg, 4Bosch Center for Artificial Intelligence, 5Columbia University, 6University of Maryland, College Park Equal contributions. bolei.ma@lmu.de, yuting.li@uni-koeln.de, wei.zhou3@de.bosch.com, zg2272@columbia.edu
@@ 1 Introduction
In linguistics, pragmatics studies how context influences the meaning of language Huang (2017); Xiang et al. (2024); Birner (2012), and how people use language in real-life situations to convey implied meanings, emotions, and intentions. Foundational work in this field, such as Grice’s (1975) work on implicature and the cooperative principle, Austin’s (1975) idea of speech acts, and Sperber and Wilson’s (1986) exploration of contextual inference laid the groundwork of the study of language use. These concepts continually influence linguistic studies and also provide insights for computational methods Mann (1980); Saygin and Cicekli (2002); Hovy and Yang (2021); Cambria (2025).
@@
However, traditional Natural Language Processing (NLP) models, while competent in syntactic parsing and semantic analysis, struggle with meaning that extends beyond the literal definition of words. This gap is where pragmatics becomes essential. Early approaches, like rule-based systems, use explicitly coded rules for knowledge representation and match input with a knowledge base for response generation Bajwa and Choudhary (2006); Grosan et al. (2011). Later, statistical models applied probability theory to predict language behavior Johnson (2009), while deep learning, with transformer-based models like BERT Devlin et al. (2019) and GPT Brown et al. (2020), transformed NLP by enabling the generation of contextually relevant text. Yet their performance in understanding pragmatics remains limited Hu et al. (2023); Sileo et al. (2022); Park et al. (2024b).
@@
Recently, the emergence of Large Language Models (LLMs) has intensified the need to evaluate their human-like communication abilities, particularly for nuanced use cases requiring sophisticated pragmatic reasoning Hu et al. (2023); Ruis et al. (2023); Yerukola et al. (2024). As LLMs are increasingly deployed in real-world applications, validating their ability to understand and generate contextually appropriate and pragmatically accurate responses is crucial to ensure effective and trustworthy human-computer interactions. While recent advances in LLMs have demonstrated impressive capabilities in generating coherent and contextually appropriate text, their pragmatic competence remains insufficiently evaluated Hu et al. (2023); Kwon et al. (2023); Chang et al. (2024).
@@
Important questions arise: What resources have been developed to evaluate the pragmatic capabilities of NLP models, and, more importantly, how can we leverage pragmatics to guide the advancement of LLMs, formalize a comprehensive framework for evaluating pragmatics, and further advance the study of pragmatics in linguistics? To answer these questions, we conduct a comprehensive survey of the resources available for evaluating pragmatics in NLP: In §2 we introduce the core concepts in pragmatics and in §3 we review how these pragmatic phenomena have been evaluated across various NLP tasks. In §4 we summarize how previous works build a pragmatic dataset. In §5 we present the metrics and evaluation techniques used in current research. In §6 we highlight gaps and offer recommendations for future research.
@@
Figure 1: A taxonomy of pragmatic phenomena.
@@ 2 Core Concepts in Pragmatics
2.1 Pragmatic Phenomena in Linguistics
@@
Context. Context is the foundation of interpretating pragmatics, encompassing both linguistic context (e.g., surrounding clauses or sentences) and extra-linguistic context, including temporal, spatial, and social factors Huang (2017).
@@
Deixis. In particular, deictic expressions rely heavily on the situational context of an utterance for their interpretation. These expressions include personal deixis (e.g., “I”, “you”), spatial deixis (e.g., “this”, “that”), and temporal deixis (e.g., “now”, “then”), and are interpreted based on factors such as the speaker’s location, the time of the utterance, and the prior discourse Levinson (2006); Stapleton (2017); Levinson (1983).
@@
Implicature. Grice (1975)’s notion of implicature describes how speakers imply additional meanings without stating them explicitly. For example, if Ann asks, “Do you sell paste?” and Bill replies, “I sell rubber cement”, the implication is that Bill does not sell paste Hirschberg (1985). Implicature results from enriching the context with additional assumptions to comply with the cooperative principle and the conversational maxims: truthfulness (Maxim of Quality), informativeness (Maxim of Quantity), relevance (Maxim of Relevance), and clarity (Maxim of Manner) (Grice, 1975).
@@
Presupposition. Presuppositions are the implicit assumptions that must hold true for an utterance to be meaningful and understood Stalnaker et al. (1977). For instance, the sentence “The Queen of England is bald” presupposes that England has a unique Queen, even though this fact is not directly stated Stalnaker et al. (1977).
@@
Speech Acts and Intent Recognition. Understanding speech acts is critical for understanding speaker intent. A speech act is an utterance that not only conveys information but also performs an action Austin (1975). Each utterance can be classified as performing an act that fulfills a communicative purpose, such as such as assertion, suggestion or description. For example, the utterance “Can you open the window?” may literally ask about the addressee’s ability or, depending on the context, function as an indirect request.
@@
Discourse and Coherence. In linguistics, discourse refers to language use beyond sentence level Guardado (2018). Discourse structure is analyzed by examining how sentences and larger text units are interconnected through coherence relations such as elaboration, explanation and contrast. Some theoretical approaches offer hierarchical models that map these relations to reveal the underlying structure of texts (for an overview of Rhetorical Structure Theory, see Mann and Thompson, 1988), while other frameworks provide a formal semantic account that integrates these coherence relations into dynamic discourse representation (for insights into Segmented Discourse Representation Theory, see Asher and Lascarides, 2003).
@@
Social Pragmatics. Social pragmatics broadens traditional theories by exploring how social factors like culture, power, gender, and interpersonal dynamics influence language use. Brown and Levinson (1987) introduced the concept of “face” (self-esteem), showing how politeness strategies manage relationships and preserve social harmony. Gender differences in communication, as examined by Tannen (1990) and Eckert and McConnell-Ginet (1992), highlight contrasting conversational styles between men and women. Cultural variations further complicate communication, as Scollon et al. (2011) demonstrate that politeness norms differ across cultures. In the digital era, social pragmatics also informs computer-mediated communication, where cues like emoticons convey emotions and tone Herring et al. (2013).
2.2 Pragmatic Phenomena in the Survey
@@
We next examine the pragmatic phenomena addressed in NLP, which were studied in the surveyed works. An overview of the survey methods and inclusivity is presented in Appendix §A. Since pragmatic concepts are interconnected, there are overlaps between categories. Therefore, we focus on the key aspects that each paper addresses. Figure 1 gives an overview of the phenomena and papers.
@@
Context and Deixis. Context and deixis in NLP evaluation tasks usually assess a model’s ability to interpret inputs based on the situational or linguistic context. The examination of the context and deixis is the basis of evaluating the models’ pragmatic ability as the models rely on the context to respond, similar to the human response process Greasley and Owen (2016); van Dijk et al. (2023); Ma et al. (2024a). Datasets or frameworks like Min et al. (2020); Sravanthi et al. (2024); Li et al. (2023a); Qi et al. (2023); Sileo et al. (2022); Hu et al. (2023); Shaikh et al. (2023); Park et al. (2024b, a); De Vries et al. (2017) explicitly address deixis as one of their pragmatic phenomena, requiring models to resolve references in context. In addition, Li et al. (2023a) and Qi et al. (2023) incorporate context-dependent reasoning, requiring models to interpret utterances within multi-turn dialogues. Monroe et al. (2017, 2018) explore language use in color reference tasks, demonstrating how context influences referential choices. Westera et al. (2020) and Zeyrek et al. (2018) also emphasize the role of context in discussion structure and question-under-discussion frameworks. These datasets highlight the role of context in grounding language understanding.
@@
Implicature and Presupposition. A recurring theme across many papers is the study of implicature, i.e., testing the model’s inference ability based on the textual input across literal meaning. It explores how meaning is implied rather than explicitly stated, requiring models to infer intentions and assumptions beyond literal interpretations. Different types of implicatures have been studied: Qi et al. (2023), Hu et al. (2023) and Halat and Atlamaz (2024) focus on conversational implicatures, testing models’ ability to infer pragmatic meanings in dialogue. Nizamani et al. (2024) investigate scalar implicatures, particularly with gradable adjectives. Zheng et al. (2021) use a grammar-based approach to generate dialogues with intricate implicatures. Cong (2024) looks into LLMs’ understanding of manner implicature, a pragmatic inference triggered by a violation of Grice (1975)’s manner maxim. Presupposition is another aspect of implicature and influences the model’s inference ability on the implications. Louis et al. (2020); Damgaard et al. (2021); Müller and Plank (2024) focus on indirectness in dialogue. Sap et al. (2020) work on implicatures posted in social media and study social biases. Kameswari et al. (2020) use pragmatic presupposition to enhance bias detection in political news. Moreover, additional works Jeretic et al. (2020); Koyano et al. (2022); Srikanth et al. (2024); Kim et al. (2024) use pragmatic inferences to understand meanings of the given input.
@@
Speech Acts and Intent Recognition. Speech acts have been explicitly studied in the surveyed works, usually in cases where there are requests, commands, or promises, given to the model and test the model’s intent recognition. Li et al. (2023a) include tasks for pragmatic identification and reasoning, requiring models to recognize speech acts in multi-turn dialogues. Reinig et al. (2024) provide a fine-grained annotation of speech acts in parliamentary debates. Braga et al. (2006) link prosodic patterns to speech acts and discourse events. A few papers have explicitly different “speakers” involved in the evaluation frameworks and assess how well the models play as the “speakers” in the dialogues, performing the commended task. Khani et al. (2018) explore how players infer intentions and generate pragmatic messages in collaborative tasks, focusing on the rational speech act between the two agents (the “speaker” and the “listener”). Shaikh et al. (2023) involve a multi-turn collaborative two-player game based on the codenames. Ollagnier (2024) also incorporates speech act-like annotations, such as “attack” and “defend”, to capture the pragmatic roles of messages in multiparty chats. Both Zhang et al. (2018) and Welleck et al. (2019) talk about the personas, who are the agents making requests, in dialogue and work on the consistency between persona and dialogue.
@@
Discourse and Coherence. Discourse and coherence have been studied in NLP with a focus on parsing text into discourse relations (e.g., Miltsakaki et al., 2004; Prasad et al., 2008, 2018; Miao et al., 2024), on discourse markers (e.g., Pandia et al., 2021; Sadlier-Brown et al., 2024), and on modeling dimensions of coherence (e.g., Lai and Tetreault, 2018; Wu et al., 2023). In addition, discourse structures play a crucial role in dialogue and communication. Westera et al. (2020) and Zeyrek et al. (2018) investigate the role of evoked questions in discourse structure using TED talks. Asher et al. (2016) explore discourse structure in multi-party dialogues and Reinig et al. (2024) study multi-party dialogues in German parliament. Moreover, discourse modes and aspect selection are influenced by contextual and pragmatic factors, connecting coherence to pragmatic interpretation (Mavridou et al., 2015). For a comprehensive overview of situation types and aspects, see Friedrich et al. (2023).
@@
Social Pragmatics. Social pragmatics explores how language use is shaped by social norms, power dynamics, and cultural contexts, including the roles of agents in communication. Ollagnier (2024) introduces a tagset for annotating discursive roles in cyberbullying, while Sap et al. (2019) and Sap et al. (2020) provide benchmarks for reasoning about social norms, biases, and power dynamics. Shaikh et al. (2023) highlight the role of shared cultural knowledge in communication through a collaborative word game, and Zhang et al. (2018) and Welleck et al. (2019) emphasize using personal and social context to improve dialogue systems. Additionally, Yang et al. (2021) examine how pragmatic features vary between neurotypical and neurodiverse speakers, particularly adults with autism. These datasets collectively demonstrate the intersection of social pragmatics with other phenomena like discourse, implicature, and speech acts.